• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name sarty

    Log entry time 07:45:41 on June 2,2000

    Entry number 42849

    keyword=shift summary

    Production running at Thcm= -135 deg. (i.e. thcm=135, phi=180)
    No problems throughout shift - other than those several messages submitted to halog by Templon regarding this shift crews lack of understanding for several spectra in hana and espace/fpp analyses.
    .
    Run 1674-1679: Production runs- no problems. (6 x 900k = 5.4M trigs)
    Run 1680: Prod run - stopped early due to rf trip followed by QE meas (330k trigs)
    Run 1681: Prod run - 900 k trigs
    Run 1682: Prod run - still running as of 08:00.
    .
    Total trigs this shift: Runs 1674-1679 = 6.63 M trigs
    .
    Total trigs (last shift and this one) on thcm=-135 deg = 6.63 M + 5.13 M = 11.76 M